OpenStudy (anonymous):

10: A bin of bolts at the hardware store contains 7-dozen bolts when full. The stock clerk is supposed to reorder bolts when the bin is 1⁄6 full. How many bolts are in the bin when it’s time to reorder?

OpenStudy (anonymous):

7 dozen is 7 x 12 which is 84 and 1/6 is rough .17 so 84 x .16 is 14

OpenStudy (anonymous):

how do i do it tho

OpenStudy (anonymous):

7 x 12 =- 84 1/6 = .17 if 84 is the total and you need 1/6 or .16 of it .... you would just divide by 6 84 / 6 = 14

OpenStudy (anonymous):

ooo so first you multiply then divide got it

OpenStudy (anonymous):

yes ..however you can also say 1/6 of 12 and then multply it by 7 because there is 7 12s in there. 1/6 of 12 is 2 x 7 = 14..is that easier?